OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (–2, 4); Slope: 5 A. y + 4 = 5x – 10 B. y + 4 = 5(x + 2) C. y – 4 = 5(x – 2) D. y – 4 = 5(x + 2)

OpenStudy (anonymous):

These kind of problems are too hard for me

OpenStudy (amonoconnor):

So, on the left with the y's there must be a 4. on the right, there must be a -2. And... M is obviously 5, so...

OpenStudy (anonymous):

also for me

OpenStudy (amonoconnor):

It's D.

OpenStudy (anonymous):

well i thought it was C.

OpenStudy (anonymous):

Which equation is graphed here? http://static.k12.com/calms_media/media/308500_309000/308804/1/ebdb4f74ba7de098e2d2acc25c822925700c2e09/MS_ALG_S1_07_08_quiz_Q5_question.gif A. y – 2 = -3/2 (x + 4) B. y – 4 = - 3/2(x + 2) C. y+4= -3/2 (x-2) D.y – 4 = -2/3 (x + 2)

OpenStudy (anonymous):

Can you help me with that last question ^

OpenStudy (amonoconnor):

Nope. The equation is: The minus and the subtraction sign on the right cancel out to make it PLUS 2.

OpenStudy (amonoconnor):

Yep, you bet! :)

OpenStudy (amonoconnor):

So we can see that the slope is: down 3, over 2 = -3/2 So... m = (-3/2) And let's use (0, -1). To find equation: y = mx + b (-1) = (-3/2)(0) + b -1 = (0) + b -1 = b Now... y = mx + b <-- general form y = (-3/2)x - 1 <-- your equation

OpenStudy (amonoconnor):

Or A. Sorry. Did it in the wrong form. Here: y - y = m(x - x) So, using (2, -4) y - 2 = (-3/2)x - (-4) y - 2 = -3/2(x + 4)

OpenStudy (anonymous):

yes thats what I was gonna pick

OpenStudy (amonoconnor):

Yay! :)

OpenStudy (anonymous):

wait dont leave because I didnt pass my test :(

OpenStudy (amonoconnor):

Okay !

OpenStudy (anonymous):

i had 2 wrong :/

OpenStudy (anonymous):

Ok I had the last question that you said it was A. it was wrong

OpenStudy (anonymous):

It was C.

OpenStudy (amonoconnor):

:(

OpenStudy (anonymous):

So im gonna retake the test can you stil help???

OpenStudy (anonymous):

I have 2 more chances

OpenStudy (amonoconnor):

Yep, and I promise to do better:)

OpenStudy (anonymous):

Ok Thanks :)

OpenStudy (anonymous):

I just have to wait a couple of minutes so my learning coach can put her password to enter the test again